The costs associated with producing six widgets are as follows: .80 variable cost per widget and .10 fixed cost per widget. Using this information, complete the following calculations:

What is the total cost of producing six widgets?

What is the average total cost of producing six widgets?

If the variable cost of producing five widgets is .60 and the fixed cost for producing five widgets is .10, what is the marginal cost of producing the sixth widget?

How would each of the following affect average total cost, average variable cost, and marginal cost?

An increase in the cost of the lease of the firm's building

A reduction in the price of electricity

A reduction in wages

A change in the salary of the president of the company

Imagine 2 other situations that would affect the costs and explain their effects.

Present your calculations and assessments of the above situations in a 2-3 page report.
Answer #1

Variable cost =. 80(6)=4.8

Fixed cost =. 10(5)=.50.This is total fixed cost which will remain same

Total cost=. 50+4.8=5.3

B average... =5.3/6=0.88

C it will include only cost of 6th unit. Here fixed cost will be.. 50/6=0.08 and variable cost is not given here. If we assume it is. 80 then marginal cost is o.88

D lease will increase only fixed cost. Average total cost will rise

E ATc, Avc and MC will decline since electricity is variable cost

F All will fall as it is variable cost

G presidency salary is fixed cost. Only ATc will fall

B per unit tax will increase all costs. Lumpsum tax is fixed cost and will raise only ATC

B